OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6

OpenStudy (anonymous):

4x-4=2x+4+6 2x=14 x=7

OpenStudy (anonymous):

4x-4=2x+4+6 4x-4=2x+10 add four to both sides 4x=2x+14 subtract 2x from both sides 2x=14 divide both sides by 2 x=7